OpenStudy (anonymous):

For each direct variation, find the constant of variation. Then find the value of y when x=-5 y=2 when x=7

OpenStudy (mayankdevnani):

for direct varaition y=kx

OpenStudy (mayankdevnani):

when y=2 and x=7

OpenStudy (mayankdevnani):

so, 2=7x

hartnn (hartnn):

*correction 2=7k

OpenStudy (mayankdevnani):

ya

OpenStudy (mayankdevnani):

so , k=2/7

OpenStudy (anonymous):

so 7/2 ? but then how do you plug in the -5

OpenStudy (mayankdevnani):

So the direct varaiation equation is y = 2/7x

OpenStudy (mayankdevnani):

when x=7 y=2/7 so k=2/7*7=2

OpenStudy (mayankdevnani):

@Ilovekevin97